# Uncovering Weaknesses in Neural Code Generation

### v1

@@ VII-A Limitations and Opportunities
Code generation that involves third-party libraries remains a significant challenge. Despite the advent of promising approaches in previous studies targeting the recommendation of third-party libraries for code generation models [60], further exploration is essential in this field. Our detailed analysis of PLMs—as discussed in Section IV-B—demonstrates that substantial difficulties persist with these issues. Moreover, the introduction of more rigorous benchmarks such as BigCodeBench [67] aims to evaluate the capability of code generation models to navigate diverse function calls and complex instructions. Concurrently, there is increasing research attention toward areas like private library recommendation [58] and unseen API recommendation [65]. We are optimistic that the emergence of such benchmarks and focused research efforts will catalyze the creation of more adept models.
Enrich benchmark datasets with answers involving multiple diverse APIs. It is widely acknowledged that current metrics, whether based on textual/semantic similarity or unit-test execution, are incapable of precisely evaluating code generation performance. Despite the continuous introduction of novel metrics such as CrystalBLEU [15] and CodeBERTScore [64], biases related to single-answer scenarios have yet to receive sufficient attention. Given the prevalence of this issue—for instance, a 43.66% accuracy rate for GPT-4 across 339 random samples from CoNaLa—we advocate for increased research efforts to establish benchmarks that offer multiple diverse API-based solutions for each prompt, enabling more robust evaluations.
VII-B Threats to Validity
@@
Manual rectification. To underscore the importance of mitigating the weaknesses in prompts, we conducted a preliminary manual correction process (See Section VI). During this phase, our modifications didn’t aim to cater specifically to the standard answer code. Instead, our focus was on promptly following the guidelines and addressing the associated weaknesses as effectively as possible [54, 37]. The modified prompts have been made publicly accessible via the artifact URL for transparency and further investigation.
@@ VIII Related Work
Recent studies have increasingly focused on systematically evaluating popular LLMs for code-related tasks, such as vulnerability detection [46], program repair [8], code summarization [44], code completion, and code generation [11, 35, 3, 61].
@@
With respect to neural code generation, numerous investigations have scrutinized the bugs in the output. For instance, Liu et al. [32] evaluate the code generated by ChatGPT from three aspects: correctness, complexity and security. Dou et al. [13] examined the performance of seven LLMs over common benchmarks, classifying bugs into syntax, runtime, and functional categories based on the Python interpreter’s feedback. Wang et al. [52] studied incorrect snippets from six LLMs on the HumanEval benchmark, constructing an error taxonomy that accounts for issues like missing conditions or incorrect return values. Tambon et al. [48] identified 10 bug patterns across 333 bugs in code generated by three LLMs, including syntax errors and wrong input types. These studies provide granular insights into the flaws within code produced by LLMs, but tend to focus solely on the errors in the generated code. Our current work aims to deliver a more holistic perspective that illuminates weaknesses in code generation, paying close attention not only to the resultant code but also to the benchmarks utilized—prompt design and reference answers included.
@@
In addition to examining the bugs, research has explored factors influencing code generation quality. Antonio et al. [35] performed an empirical study showing that variances in natural language descriptions could affect the output of GitHub Copilot. Cao et al. [7] observed similar impacts across different datasets and models. Jin et al. [22] found that carefully crafted prompts significantly improve zero-shot performance, and Lu et al. [33] demonstrated the sensitivity of prompt ordering to few-shot learning success, devising a method to identify effective prompts without additional annotated data.
@@
Shi et al. [44] emphasized the role of data preprocessing in the performance of state-of-the-art code summarization models. Zan et al. [61] identified three critical factors for LLM success in the NL2Code task: model size, high-quality data, and expert tuning. Liu et al. [31] pointed out test inadequacy and vague prompt descriptions in benchmarks, proposing EvalPlus, a framework capable of generating a broader set of test cases to expose more nuanced issues. Moradi Dakhel et al. [38] assessed Copilot against human programmers, noting its shortcomings in synthesizing complex solutions and its occasional neglect of key prompt details—a phenomenon our research categorizes as Type V: Missing pivotal semantics.
@@
The aforementioned studies have also provided crucial insights for improving related models. For instance, Seif et al. [3] discovered that UML-enhanced prompts boost the reliability of generated Object Constraint Language (OCL) outputs. Other research has demonstrated the benefits of including context elements, like variable names and function arguments, to refine code generation [1][21]. The impact of external knowledge has been similarly affirmed: Liu et al. [30] showed that it enhances commonsense reasoning in sequence models, suggesting that prompting with domain-specific information (Type VII) might improve code generation.
@@
Some researchers have also highlighted benchmark limitations. Liu et al. [28] warned that models potentially overfit on well-trodden datasets like HumanEval, calling for more diverse datasets for authentic assessment of code generation. Huang et al. [20] introduced EffiBench, a new benchmark focusing on the efficiency of model-generated code, urging consideration of not just correctness but also the efficiency and sustainability of code outputs.
IX Conclusions and Future work

### v2

@@ 7.1. Limitations and Opportunities
  • Code generation that involves third-party libraries remains a significant challenge. Despite the advent of promising approaches in previous studies targeting the recommendation of third-party libraries for code generation models (Zan et al., 2022c), further exploration is essential in this field. Our detailed analysis of PLMs—as discussed in Section 4.2—demonstrates that substantial difficulties persist with these issues. Moreover, the introduction of more rigorous benchmarks such as BigCodeBench (Zhuo et al., 2024) aims to evaluate the capability of code generation models to navigate diverse function calls and complex instructions. Concurrently, there is increasing research attention toward areas like private library recommendation (Zan et al., 2022a) and unseen API recommendation (Zhou et al., 2023b). We are optimistic that the emergence of such benchmarks and focused research efforts will catalyze the creation of more adept models.
  • Enrich benchmark datasets with answers involving multiple diverse APIs. It is widely acknowledged that current metrics, whether based on textual/semantic similarity or unit-test execution, are incapable of precisely evaluating code generation performance. Despite the continuous introduction of novel metrics such as CrystalBLEU (Eghbali and Pradel, 2022) and CodeBERTScore (Zhou et al., 2023a), biases related to single-answer scenarios have yet to receive sufficient attention. Given the prevalence of this issue—for instance, a 43.66% accuracy rate for GPT-4 across 339 random samples from CoNaLa—we advocate for increased research efforts to establish benchmarks that offer multiple diverse API-based solutions for each prompt, enabling more robust evaluations.
7.2. Threats to Validity
@@
  • Manual rectification. To underscore the importance of mitigating the weaknesses in prompts, we conducted a preliminary manual correction process (See Section 6). During this phase, our modifications didn’t aim to cater specifically to the standard answer code. Instead, our focus was on promptly following the guidelines and addressing the associated weaknesses as effectively as possible (White et al., 2023; Mishra et al., 2022). The modified prompts have been made publicly accessible via the artifact URL for transparency and further investigation.
@@ 8. Related Work
Recent studies have increasingly focused on systematically evaluating popular LLMs for code-related tasks, such as vulnerability detection (Steenhoek et al., 2023), program repair (Cao et al., 2023), code summarization (Shi et al., 2022), code completion, and code generation (Ciniselli et al., 2021; Mastropaolo et al., 2023; Abukhalaf et al., 2023; Zan et al., 2023).
@@
With respect to neural code generation, numerous investigations have scrutinized the bugs in the output. For instance, Liu et al. (Liu et al., 2024a) evaluate the code generated by ChatGPT from three aspects: correctness, complexity and security. Dou et al. (Dou et al., 2024) examined the performance of seven LLMs over common benchmarks, classifying bugs into syntax, runtime, and functional categories based on the Python interpreter’s feedback. Wang et al. (Wang et al., 2024) studied incorrect snippets from six LLMs on the HumanEval benchmark, constructing an error taxonomy that accounts for issues like missing conditions or incorrect return values. Tambon et al. (Tambon et al., 2024) identified 10 bug patterns across 333 bugs in code generated by three LLMs, including syntax errors and wrong input types. These studies provide granular insights into the flaws within code produced by LLMs, but tend to focus solely on the errors in the generated code. Our current work aims to deliver a more holistic perspective that illuminates weaknesses in code generation, paying close attention not only to the resultant code but also to the benchmarks utilized—prompt design and reference answers included.
@@
In addition to examining the bugs, research has explored factors influencing code generation quality. Antonio et al. (Mastropaolo et al., 2023) performed an empirical study showing that variances in natural language descriptions could affect the output of GitHub Copilot. Cao et al. (Cao et al., 2022) observed similar impacts across different datasets and models. Jin et al. (Jin et al., 2022) found that carefully crafted prompts significantly improve zero-shot performance, and Lu et al. (Lu et al., 2022) demonstrated the sensitivity of prompt ordering to few-shot learning success, devising a method to identify effective prompts without additional annotated data.
@@
Shi et al. (Shi et al., 2022) emphasized the role of data preprocessing in the performance of state-of-the-art code summarization models. Zan et al. (Zan et al., 2023) identified three critical factors for LLM success in the NL2Code task: model size, high-quality data, and expert tuning. Liu et al. (Liu et al., 2023) pointed out test inadequacy and vague prompt descriptions in benchmarks, proposing EvalPlus, a framework capable of generating a broader set of test cases to expose more nuanced issues. Moradi Dakhel et al. (Moradi Dakhel et al., 2023) assessed Copilot against human programmers, noting its shortcomings in synthesizing complex solutions and its occasional neglect of key prompt details—a phenomenon our research categorizes as Type V: Missing pivotal semantics.
@@
The aforementioned studies have also provided crucial insights for improving related models. For instance, Seif et al. (Abukhalaf et al., 2023) discovered that UML-enhanced prompts boost the reliability of generated Object Constraint Language (OCL) outputs. Other research has demonstrated the benefits of including context elements, like variable names and function arguments, to refine code generation (Con, [n. d.])(Iyer et al., 2018). The impact of external knowledge has been similarly affirmed: Liu et al. (Liu et al., 2022) showed that it enhances commonsense reasoning in sequence models, suggesting that prompting with domain-specific information (Type VII) might improve code generation.
@@
Some researchers have also highlighted benchmark limitations. Liu et al. (Liu et al., 2024b) warned that models potentially overfit on well-trodden datasets like HumanEval, calling for more diverse datasets for authentic assessment of code generation. Huang et al. (Huang et al., 2024) introduced EffiBench, a new benchmark focusing on the efficiency of model-generated code, urging consideration of not just correctness but also the efficiency and sustainability of code outputs.
9. Conclusions and Future work
@@
Acknowledgments
We extend our gratitude to the additional participants who contributed to data annotation beyond the authors. Our thanks go to Feng Luo, Zixuan Zhai, and Jiehua Li from Harbin Institute of Technology, as well as Jiale Xu, ChengHu, Ziyan Zhao, Donghao Yang, Tiansheng Lin, and Shaoming Zhu, from Beihang University. Funding for this work has been provided by the National Science Foundation of China Grant No.62102014 and supported by State Key Laboratory of Software Development Environment No.SKLSDE-2023ZX-10.